# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 4.2 Convergence to stationary points
$\leq$ $\|\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\|+(L_{f_{x}}+L_{g_{2}})C_{\eta}s_{k},$
where the last inequality follows from (35). Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and assuming that $\nabla_{x}F,\nabla_{x}f$ and $\nabla_{x}g$ are continuous on $X\times Y$, and $\nabla_{x}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we conclude that $\nabla_{x}\tilde{\psi}_{p_{k}}=\nabla_{x}\tilde{\psi}_{\bar{p}}$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}})=(\bar{x},\bar{y})$ and $s_{k}\rightarrow 0$, it follows that $\nabla_{x}\tilde{\psi}_{p_{k_{j}}}(x^{k_{j}+1},y^{k_{j}+1})-d_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with (40) and (43) and the fact that $\alpha_{k}\geq\underline{\alpha}>0$, we obtain that $\xi_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. A similar argument can be applied to $\xi_{y}^{k}$. From (30), we have
@@
$\leq$ $\|\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k}\right)\|+{C_{\eta}}s_{k}/\gamma.$
Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and under the assumption that $\nabla_{y}F$ and $\nabla_{y}f$ are continuous on $X\times Y$, and $\nabla_{y}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we have that $\nabla_{y}(\tilde{\psi}_{p_{k}}-g)=\nabla_{y}(\tilde{\psi}_{\bar{p}}-g)$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}})=(\bar{x},\bar{y})$ and $s_{k}\rightarrow 0$, it follows that $\nabla_{y}(\tilde{\psi}_{p_{k_{j}}}-g)(x^{k_{j}},y^{k_{j}+1})-d_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with (40), (45) and the fact that $\beta_{k}\geq\underline{\beta}>0$, we conclude that $\xi_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
Now, taking the limit as $k=k_{j}\rightarrow\infty$ in (42), since $\xi_{x}^{k_{j}},\xi_{y}^{k_{j}}\rightarrow 0$, $p_{k}=\bar{p}$ for all $k\geq k_{0}$, $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}})=(\bar{x},\bar{y})$, $\nabla F$, $\nabla v_{\gamma}$, and $\nabla_{x}\varphi$ are continuous on $X\times Y$, $\partial_{y}\varphi$ and $\mathcal{N}_{X\times Y}$ are outer semicontinuous on $X\times Y$, and $\partial\varphi=\{\nabla_{x}\varphi\}\times\partial_{y}\varphi$ on $X\times Y$, we obtain
@@
Finally, since $s_{k}\rightarrow 0$, we have $\|\theta^{k}-\theta^{*}(x^{k},y^{k})\|\rightarrow 0$, and thus $\varphi(x^{k_{j}},\theta^{k_{j}})-\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$ from the continuity of $\varphi$ and $\theta^{*}(x,y)$ (Proposition 2). Taking the limit as $k=k_{j}\rightarrow\infty$ in (41), we have
$\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon.$ (47)
Therefore, letting $\bar{\epsilon}:=\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})$, we have $\bar{\epsilon}\in[0,\epsilon]$ and together with (45), we obtain
@@
As established in Lemma 6, the Ostrowski condition $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\|=0$ is satisfied. By combining this result with Theorem 2 and [55, Proposition 8.3.10], we derive the following convergence result:
@@
Let $\epsilon>0$. Suppose the sequence ${p_{k}}$ is bounded, and the sequence ${(x^{k},y^{k})}$ has an isolated accumulation point $(\bar{x},\bar{y})$, meaning there exists a neighborhood $Z$ of $(\bar{x},\bar{y})$ such that $(\bar{x},\bar{y})$ is the only accumulation point of ${(x^{k},y^{k})}$ in $Z$. Then the sequence ${(x^{k},y^{k})}$ converges to $(\bar{x},\bar{y})$, which is a KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
As demonstrated in Proposition 3, ENNAMCQ holds for $({\rm VP})_{\gamma}^{\epsilon}$ for any $\epsilon>0$ when the lower level problem is strictly convex with respect to the lower level variable $y$. Next, we show that ENNAMCQ is sufficient to guarantee the boundedness of the penalty parameter $p_{k}$ generated by Algorithm 1, as required in Theorem 2.
Theorem 4.
Let $\epsilon>0$. If the sequence $\left\{(x^{k},y^{k})\right\}$ is bounded, and ENNAMCQ holds at any accumulation points of $\left\{x^{k},y^{k}\right\}$, then $\left\{p_{k}\right\}$ must be bounded.
Proof.
The proof builds upon techniques used in the proof of Proposition 3 in [27]. We proceed by contradiction, assuming that $p_{k}\to\infty$ as $k\to\infty$. First, recall from Lemma 6 that $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\|=0$. From the update rule of $p_{k}$, there exist a subsequence $\{k_{j}\}\subseteq\mathbb{N}$ such that
$t^{k_{j}+1}>c_{p}^{-1}\left\|\left(x^{k_{j}},y^{k_{j}}\right)-\left(x^{k_{j}+1},y^{k_{j}+1}\right)\right\|>0.$ (49)
Since the sequence $\{(x^{k},y^{k})\}$ is bounded, we can assume without loss of generality by extracting a further subsequence that $(x^{k_{j}},y^{k_{j}})\to(\bar{x},\bar{y})$ for some $(\bar{x},\bar{y})\in X\times Y$. Suppose $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$. As shown in the proof of Theorem 2, $\varphi(x^{k_{j}},\theta^{k_{j}})+\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. Thus, there exists $k_{j}$ such that
$\varphi(x^{k_{j}},y^{k_{j}})-\varphi(x^{k_{j}},\theta^{k_{j}})-\frac{1}{2\gamma}\left\|\theta^{k_{j}}-y^{k_{j}}\right\|^{2}-\epsilon<0,$
which implies $t^{k_{j}}=0$. This contradicts (49). Therefore, we conclude that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon\geq 0$. Next, consider(42), which states
$p_{k}^{-1}\left(\xi_{x}^{k}-\nabla F(x^{k+1},y^{k+1})\right)$ $\in\nabla_{x}\varphi(x^{k+1},y^{k+1})-\nabla_{x}v_{\gamma}(x^{k+1},y^{k+1})+\mathcal{N}_{X}(x^{k+1}),$ (50)
$p_{k}^{-1}\left(\xi_{y}^{k}-\nabla F(x^{k},y^{k+1})\right)$ $\in\partial_{y}\varphi_{y}(x^{k},y^{k+1})-\nabla_{y}v_{\gamma}(x^{k},y^{k+1})+\mathcal{N}_{Y}(y^{k+1})$
where $\xi_{x}^{k},\ \xi_{y}^{k}$ are defined in (43). Using similar arguments as in the proof of Theorem 2, we can show that
$\lim_{k_{j}\to\infty}\left(p_{k_{j}}^{-1}\left(\xi_{x}^{k_{j}}-\nabla_{x}F(x^{{k_{j}}+1},y^{{k_{j}}+1})\right),p_{k_{j}}^{-1}\left(\xi_{y}^{k_{j}}-\nabla_{y}F(x^{{k_{j}}},y^{{k_{j}}+1})\right)\right)=0.$
Taking the limit as $k=k_{j}\rightarrow\infty$ in (50) yields
$0\in\nabla f(\bar{x},\bar{y})+\partial g(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}),$
which contradicts the assumption that ENNAMCQ holds and thus we get the conclusion. ∎
@@ 5 Sequential convergence under KL property
This section establishes the sequential convergence of AGILS (Algorithm 1) under the Kurdyka-Łojasiewicz (KL) property. We begin by recalling the definition of the KL property. Let $\zeta\in[0,+\infty)$ and $\Phi_{\zeta}$ represent the set of all concave and continuous functions $\phi\ :\ [0,\zeta)\to[0,+\infty)$ that meet the following conditions: (a) $\phi(0)=0$, (b) $\phi$ is $C^{1}$ on $(0,\zeta)$ and continuous at 0, (c) $\phi^{\prime}(s)>0$ for all $s\in(0,\zeta)$.
@@
Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (Algorithm 1) under the step size conditions specified in Section 4.2, where $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants. To analyze sequential convergence, we define the following merit function $E_{p}(z,u)$, where $z:=(x,y)$,
@@
$G_{p}(z):=\frac{1}{p}F(z)+\varphi(z)+\frac{\rho_{v_{1}}}{2}\|x\|^{2},\quad H(u,y):=\sup_{x}\{\langle u,x\rangle-\tilde{v}_{\gamma}(x,y)\},$
where $\rho_{v_{1}}=\rho_{f_{1}}+\rho_{g_{1}}+c_{\alpha}/4$ and $\tilde{v}_{\gamma}(x,y):={v}_{\gamma}(x,y)+(\rho_{v_{1}}/2)\|x\|^{2}$. By Proposition 2 and under Assumption 3, $\tilde{v}_{\gamma}(x,y)$ is strongly convex with respect to $x$. $H(x,u)$ is the Fenchel conjugate of $\tilde{v}_{\gamma}(x,y)$ with respect to variable $x$ and is differentiable due to the differentiability of ${v}_{\gamma}(x,y)$ and the strong convexity of ${v}_{\gamma}(x,y)$ with respect to $x$. For the sequence $\{(x^{k},y^{k})\}$ generated by AGILS(1), let $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. We have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$ and $\nabla_{y}H(u^{k},y^{k+1})=\nabla_{y}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. The following lemma establishes a relaxed sufficient decrease property and relative error condition for the merit function $E_{p}(z,u)$ at $(z^{k+1},u^{k})$.
@@
Under Assumptions 1, 2, 3 and 4. Let $\epsilon>0$. Let $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ be the sequences generated by AGILS under the same conditions as in Section 4.2, and let $E_{p}(z,u)$ be defined as in (51), then there exist $a,b>0$ such that
@@
where $\nu_{k}=\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$, and
@@
where $\tilde{\nu}_{k}=(L_{f_{x}}+L_{g_{2}}+1/\gamma)C_{\eta}s_{k}$.
@@
We begin by deriving the first inequality. Using (31), it follows that:
$G_{p_{k}}(x^{k},y^{k+1})-{v}_{\gamma}(x^{k},y^{k+1})$ (54)
$\leq$ $G_{p_{k}}(x^{k},y^{k})-{v}_{\gamma}(x^{k},y^{k})-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}s_{k}^{2}.$
Next, using (32) and the fact that the gradients $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$ and $\nabla_{x}g(x,y^{k+1})$ are $L_{F_{x}}$, $L_{f_{x}}$ and $L_{g_{1}}$-Lipschitz continuous with respect to variable $x$ on $X$, respectively, we have
@@
Using the definitions of $G_{p}(z)$ and $\tilde{v}_{\gamma}(z)$, along with the gradient formulas for $\nabla_{x}{v}_{\gamma}(z)$ and $d_{x}^{k}$ in (19), we further derive
$G_{p_{k}}(z^{k+1})-\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k+1}-x^{k}\rangle+\left(\frac{1}{\alpha_{k}}-\frac{L_{E_{x,k}}}{2}\right)\left\|x^{k+1}-x^{k}\right\|^{2}$ (56)
$\leq$ $G_{p_{k}}(x^{k},y^{k+1})+\left\langle\nabla_{x}\varphi(x^{k},\theta^{k+1/2})-\nabla_{x}v_{\gamma}(x^{k},y^{k+1}),x^{k+1}-x^{k}\right\rangle,$
where $L_{E_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{v_{1}}$. Combining (35) with (56) and applying the Cauchy-Schwarz inequality, we obtain
@@
$\leq$ $G_{p_{k}}(x^{k},y^{k+1})+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}s_{k}^{2}.$
Combining this result with (54), we have
@@
$\leq$ $G_{p_{k}}(z^{k})-\tilde{v}_{\gamma}(z^{k})+\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$
@@
We have $-\tilde{v}_{\gamma}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k}\rangle=H(u^{k},y^{k+1})$, because $\tilde{v}_{\gamma}$ is convex in $x$ and $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. Additionally, since $-\tilde{v}_{\gamma}(x^{k},y^{k})\leq-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$, substituting this into (58) yields
@@
where $\nu_{k}=\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$, and (52) follows.

### v2

@@ 4.3 Convergence to stationary points
$\leq$ $\|\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\|+(L_{f_{x}}+L_{g_{2}})C_{\eta}\zeta_{k},$
where the last inequality follows from (39). Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and assuming that $\nabla_{x}F,\nabla_{x}f$ and $\nabla_{x}g$ are continuous on $X\times Y$, and $\nabla_{x}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we conclude that $\nabla_{x}\tilde{\psi}_{p_{k}}=\nabla_{x}\tilde{\psi}_{\bar{p}}$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$ and $\zeta_{k}\rightarrow 0$, it follows that $\nabla_{x}\tilde{\psi}_{p_{k_{j}}}(x^{k_{j}+1},y^{k_{j}+1})-d_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with $\|x^{k_{j}+1}-x^{k_{j}}\|\to 0$ and $\alpha_{k}\geq\underline{\alpha}>0$, we obtain from (47) that $\xi_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
A similar argument can be applied to $\xi_{y}^{k}$. From (34), we have
@@
$\leq$ $\|\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},\tilde{y}^{k}\right)\|+{C_{\eta}}\zeta_{k}/\gamma.$
Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and under the assumption that $\nabla_{y}F$ and $\nabla_{y}f$ are continuous on $X\times Y$, and $\nabla_{y}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we have that $\nabla_{y}(\tilde{\psi}_{p_{k}}-g)=\nabla_{y}(\tilde{\psi}_{\bar{p}}-g)$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$ and $\zeta_{k}\rightarrow 0$, it follows that $\nabla_{y}(\tilde{\psi}_{p_{k_{j}}}-g)(x^{k_{j}},y^{k_{j}+1})-d_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with $\|y^{k_{j}+1}-\tilde{y}^{k_{j}}\|\to 0$ and $\beta_{k}\geq\underline{\beta}>0$, we conclude from (49) that $\xi_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
Now, taking the limit as $k=k_{j}\rightarrow\infty$ in (46), since $\xi_{x}^{k_{j}},\xi_{y}^{k_{j}}\rightarrow 0$, $p_{k}=\bar{p}$ for all $k\geq k_{0}$, $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$, $\nabla F$, $\nabla v_{\gamma}$, and $\nabla_{x}\varphi$ are continuous on $X\times Y$, $\partial_{y}\varphi$ and $\mathcal{N}_{X\times Y}$ are outer semicontinuous on $X\times Y$, and $\partial\varphi=\{\nabla_{x}\varphi\}\times\partial_{y}\varphi$ on $X\times Y$, we obtain
@@
Finally, from Proposition 3, we know that the limit point $(\bar{x},\bar{y})$ satisfies $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$. Let $\bar{\epsilon}:=\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\in[0,\epsilon]$. Combining this with (50), we have
@@
As established in Lemma 6, the Ostrowski condition $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$ is satisfied. By combining this result with Theorem 2 and [25, Proposition 8.3.10], we derive the following convergence result:
@@
Let $\epsilon>0$. Suppose the sequence ${p_{k}}$ is bounded, and the sequence $\{(x^{k},y^{k})\}$ has an isolated accumulation point $(\bar{x},\bar{y})$, meaning there exists a neighborhood $Z$ of $(\bar{x},\bar{y})$ such that $(\bar{x},\bar{y})$ is the only accumulation point of $\{(x^{k},y^{k})\}$ in $Z$. Then the sequence $\{(x^{k},y^{k})\}$ converges to $(\bar{x},\bar{y})$, which is a KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
Remark 4.
Our algorithm AGILS offers significant advantages over the MEHA in [42], particularly regarding step-size selection and convergence guarantees. First, AGILS allows a larger, explicit, and easily computable range of step sizes. In contrast, the step sizes for MEHA must be small and lack a clear, practical selection range. This makes AGILS more flexible and easier to tune in practice. Second, AGILS ensures that any accumulation point of the iterates is a feasible KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$. Conversely, MEHA’s convergence analysis only ensures that the stationary residual of penalized problem (12) converges to zero, without guaranteeing the feasibility of the iterates.
@@ 5 Sequential convergence under KL property
This section establishes the sequential convergence of AGILS (Algorithm 1) under the Kurdyka-Łojasiewicz (KL) property. Our analysis focuses on the scenario where the feasibility correction procedure is applied only a finite number of times. We begin by recalling the definition of the KL property. Let $\zeta\in[0,+\infty)$ and $\Phi_{\zeta}$ represent the set of all concave and continuous functions $\phi\ :\ [0,\zeta)\to[0,+\infty)$ that meet the following conditions: (a) $\phi(0)=0$, (b) $\phi$ is $C^{1}$ on $(0,\zeta)$ and continuous at 0, (c) $\phi^{\prime}(s)>0$ for all $s\in(0,\zeta)$.
@@
Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (Algorithm 1) under the step size conditions specified in Section 4.2 and 4.3, where $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants. To analyze sequential convergence, we define the following merit function $E_{p}(z,u)$, where $z:=(x,y)$,
@@
$G_{p}(z):=\frac{1}{p}\left(F(z)-\underline{F}\right)+\varphi(z)+\frac{\rho_{v_{1}}}{2}\|x\|^{2},\quad H(u,y):=\sup_{x}\{\langle u,x\rangle-\tilde{v}_{\gamma}(x,y)\},$
where $\rho_{v_{1}}=\rho_{f_{1}}+\rho_{g_{1}}+c_{\alpha}/4$ and $\tilde{v}_{\gamma}(x,y):={v}_{\gamma}(x,y)+(\rho_{v_{1}}/2)\|x\|^{2}$. By Proposition 2 and under Assumption 3, $\tilde{v}_{\gamma}(x,y)$ is strongly convex with respect to $x$. $H(u,y)$ is the Fenchel conjugate of $\tilde{v}_{\gamma}(x,y)$ with respect to variable $x$ and is differentiable due to the differentiability of ${v}_{\gamma}(x,y)$ and the strong convexity of ${v}_{\gamma}(x,y)$ with respect to $x$. For the sequence $\{(x^{k},y^{k})\}$ generated by AGILS(Algorithm 1), let $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. We have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$ and $\nabla_{y}H(u^{k},y^{k+1})=\nabla_{y}\tilde{v}_{\gamma}(x^{k},y^{k+1})$.
The following lemma establishes a relaxed sufficient decrease property and relative error condition for the merit function $E_{p}(z,u)$ at $(z^{k+1},u^{k})$.
@@
Under Assumptions 1, 2, 3 and 4. Let $\epsilon>0$. Let $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ be the sequences generated by AGILS under the same conditions as in Section 4.2, and let $E_{p}(z,u)$ be defined as in (52). In the case where $\tilde{y}^{k}=y^{k}$, there exist constants $a,b>0$ such that
@@
where $\nu_{k}=\left(\frac{(L_{f_{x}}+L_{g_{2}})^{2}}{c_{\alpha}}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}\zeta_{k}^{2}$ with $\zeta_{k}:=\max\{s_{k},(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})\}$, and
@@
where $\tilde{\nu}_{k}=(L_{f_{x}}+L_{g_{2}}+1/\gamma)C_{\eta}\zeta_{k}$.
@@
We begin by deriving the first inequality. In this case, $\tilde{y}^{k}=y^{k}$, the inequality (35) simplifies to $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)-\left(1/\beta_{k}-L_{\psi_{y,k}}/2-c_{\beta}/4\right)\left\|y^{k+1}-y^{k}\right\|^{2}+C_{\eta}^{2}/(c_{\beta}\gamma^{2})\zeta_{k}^{2}$, which implies
$G_{p_{k}}(x^{k},y^{k+1})-{\tilde{v}}_{\gamma}(x^{k},y^{k+1})$ (55)
$\leq$ $G_{p_{k}}(x^{k},{y}^{k})-{\tilde{v}}_{\gamma}(x^{k},{y}^{k})-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}\zeta_{k}^{2}.$
Next, by Lemma 5, $\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k},\,x^{k+1}-x^{k}\rangle-\left(1/\alpha_{k}-L_{\psi_{x,k}}/2\right)\|x^{k+1}-x^{k}\|^{2}$. Moreover, since the gradients $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$, and $\nabla_{x}g(x,y^{k+1})$ are $L_{F_{x}}$, $L_{f_{x}}$ and $L_{g_{1}}$-Lipschitz continuous with respect to variable $x$ on $X$, respectively, we have
@@
Using the definitions of $G_{p}(z)$ and $\tilde{v}_{\gamma}(z)$, along with the formulas for $\nabla_{x}{v}_{\gamma}(z)$ and $d_{x}^{k}$, as well as (39) and applying the Cauchy-Schwarz inequality, we further derive
@@
$\leq$ $G_{p_{k}}(x^{k},y^{k+1})+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}\zeta_{k}^{2},$
where $L_{E_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{v_{1}}$. Combining this result with (55), we have
@@
$\leq$ $G_{p_{k}}(z^{k})-\tilde{v}_{\gamma}(z^{k})+\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}\zeta_{k}^{2}$
@@
By the convexity of $\tilde{v}_{\gamma}$ with respect to $x$ and $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have the identity $-\tilde{v}_{\gamma}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k}\rangle=H(u^{k},y^{k+1})$. Thus, the left-hand side of inequality (58) equals $G_{p_{k}}(z^{k+1})-\langle u^{k},x^{k+1}\rangle+H(u^{k},y^{k+1})$. Furthermore, since $-\tilde{v}_{\gamma}(x^{k},y^{k})\leq-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$, the term $G_{p_{k}}(z^{k})-\tilde{v}_{\gamma}(z^{k})$ on the right-hand side of (58) equals $G_{p_{k}}(z^{k})-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$.
Substituting these into (58) and recalling the definition of the merit function $E_{p_{k}}(z,u)$ directly yields the following inequality
$E_{p_{k}}(z^{k+1},u^{k})-E_{p_{k}}(z^{k},u^{k-1})-\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}\zeta_{k}^{2}$
$\leq$ $\left(\frac{1}{\alpha_{k}}-\frac{L_{E_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}+\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}.$
[…]
$\frac{1}{\alpha_{k}}-\frac{L_{E_{x,k}}}{2}-\frac{c_{\alpha}}{4}\geq\frac{L_{\psi_{x,k}}}{2}+\frac{c_{\alpha}}{2}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{8}-\frac{c_{\alpha}}{4}=\frac{c_{\alpha}}{8}.$
[…]
$\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\geq\frac{L_{\psi_{y,k}}}{2}+\frac{c_{\beta}}{2}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}=\frac{c_{\beta}}{4}.$
Inserting these bounds into the previous descent estimate yields the inequality
@@
where $\nu_{k}=\left(\frac{(L_{f_{x}}+L_{g_{2}})^{2}}{c_{\alpha}}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}\zeta_{k}^{2}$, and (53) follows.